The grades in a statistics course for a particular semester were as follows:

x= 1, 2, 3, 4, 5, 6, 7, 8

f= 136, 60, 34, 12, 55, 1, 3, 1

Test the hypothesis, at the 0.05 level of significance, does the distribution of grades

follow poisson distribution.
1
Expert's answer
2021-09-28T20:37:58-0400

Let X denote the grades in a statistics course.

Hypothesis tested is,

H0: X ~ Poisson distribution against H1: X does not follow a Poisson distribution

We first compute the parameter "\lambda=(\displaystyle\sum{(f*x)})/n=716/302=2.37(2 decimal places)."

We then compute respective probabilities for "x=1,2,3,4,5,6,7,8" using the Poisson distribution given by;

"P"x "=P(X=x)=(e^{-\lambda}*\lambda^x)/x!"

Therefore,

P1=P(X=1)=e-2.37*(2.37)=0.22

P2=P(X=2)=e-2.37*(2.372)/2!=0.26

P3=P(X=3)=e-2.37*(2.373)/3!=0.21

P4=P(X=4)=e-2.37*(2.374)/4!=0.12

P5=P(X=5)=e-2.37*(2.375)/5!=0.06

P6=P(X=6)=e-2.37*(2.376)/6!=0.02

P7=P(X=7)=e-2.37*(2.377)/7!=0.008

P8=P(X=8)=e-2.37*(2.378)/8!=0.002

Next is to determine the expected frequencies, EX=n*PX

E1=n*p1=302*0.22=66.87

E2=n*p2=302*0.26=79.28

E3=n*p3=302*0.21=62.65

E4=n*p4=302*0.12=37.13

E5=n*p5=302*0.06=17.61

E6=n*p6=302*0.02=6.96

E7=n*p7=302*0.008=2.36

E8=n*p8=302*0.02=0.70

Since the expected frequencies for x=7 and x=8 are less than 5, they are both combined with the expected frequency for x=6 to get a total expected frequency of 10.01.

With these values, we can now find the chi square test statistic defined by,

"\chi"2(calculated)="\displaystyle\sum_{i=1}^{6}(O_{i}-E_{i})^2/E_i"

=(136-66.87)2/66.87+(60-79.28)2/79.28+(34-62.65)2/62.65+(12-37.13)2/37.13+(55-17.62)2/17.62+(5-10.01)2/10.01=188.17(2 decimal places).

"\chi"2  table value with k-1= 6-1=5degrees of freedom at "\alpha" =0.05=11.0705.

"\chi^2"(calculated)=188.17 is greater than the table value of 11.0705 hence, we reject the null hypothesis that X follows a Poisson distribution and conclude that there is no sufficient evidence to show that grades in a statistics class follow a Poisson distribution.